I need help figuring out how to do three y squared plus five y equals twelve? I am really confused and all your help will be appreciated!
Answer by flame8855(424)

3Y^2 +5Y=12
3Y^2+5Y-12=0
(3Y-4)(Y+3)=0
Y=4/3
Y=-3